OpenStudy (anonymous):

Write the equation of the line with a slope of -5 and a y-intercept of (0,3). Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). Write the equation of the line passing through the points (0, -4) and (-2, 2). Write the equation of the line passing through the points (-6,1) and (-4,2). Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (anonymous):

im so stuck

OpenStudy (anonymous):

can anyone help please

OpenStudy (texaschic101):

slope of -5 and y intercept of (0,3) y = -5x + 3 =============== point slope formula : y - y1 = m(x - x1) slope(m) = -1/3 (6,-4) x1 = 6 and y1 = -4 sub it in y - (-4) = -1/3(x - 6) y + 4 = -1/3x + 2 y = -1/3x + 2 - 4 y = -1/3x -2 ================ slope formula : m = (y2 - y1) / (x2 - x1) (0,-4) x1 = 0 and y1 = -4 (-2,2) x1 = -2 and y1 = 2 sub it in m = (2 - (-4) / (-2 - 0) m = (2 + 4) / -2 m = 6/-2 m = -3 point slope formula : y - y1 = m(x - x1) slope(m) = -3 (-2,2) x1 = -2 and y1 = 2 sub it in y - 2 = -3(x - (-2) y - 2 = -3(x + 2) y - 2 = -3x - 6 y = -3x - 6 + 2 y = -3x - 4 ================= slope formula : m = (y2 - y1) / (x2 - x1) (-6,1) x1 = -6 and y1 = 1 (-4,2) x2 = -4 and y2 = 2 sub it in m = (2 - 1) / (-4,- (-6) m = 1/ (-4 + 6) m = 1/2 y - y1 = m(x - x1) slope(m) = 1/2 (-6,1) x1 = -6 and y1 = 1 sub it in y - 1 = 1/2(x - (-6) y - 1 = 1/2(x + 6) y - 1 = 1/2x + 3 y = 1/2x + 3 + 1 y = 1/2x + 4 ================== Undefined slope means that the slope is divided by 0. There is no " run ". It is a vertical line. If it passes through (2,5), then the equation is x = 2 ================== any questions ?